OpenStudy (xventurientalex):

What is the solution to the equation? a + (–0.4) = –3 I WILL FAN AND MEDAL TO BEST EXPLANATION AND AWNSER

OpenStudy (anonymous):

there is a math tab on here for people to help you with your math just go to find more subjects :)

OpenStudy (xventurientalex):

you dont really help much :) btw they dont awnser at all:(

OpenStudy (anonymous):

Simplify brackets a+−0.4=−3 Simplify a+−0.4 to a−0.4 a−0.4=−3 Add 0.4 to both sides a=−3+0.4 Simplify −3+0.4 to −2.6 a=−2.6

OpenStudy (anonymous):

Are you serious. Okay. well the answer is a= -2.6

OpenStudy (xventurientalex):

2.6 cause all i seeare a bunch of these ????????????????????????????????????????????????????????

OpenStudy (anonymous):

I showed the steps and how to do it but it didn't work. Give me a second. http://www.cymath.com/answer.php?q=a%2B%28-0.4%29%3D-3

OpenStudy (xventurientalex):

cool thx